In a vacuum of 10¬3 Torr at room temperature (20 °C). This number is a useful one for the experimentalist to know by heart (10¬³ Torr=1 µ).


1
Expert's answer
2021-09-18T14:58:50-0400

We know that


"PV=nkT,\\\space\\
N=\frac NV=\frac{P}{kT}=\frac{10^{-3}·133}{1.38·10^{-23}·(273+20)}=3.29·10^{-19}\text{ m}^{-3}."